Canada produces about 1.5 million tonnes of chlorine per year by the electrolysis of brine (NaCl(aq)). If the required electrolytic cells are operated at a potential difference of 4.5 V, calculate how many total kilowatt hours (kWh) of electrical power are required. 1 kWh = 3.6 × 106 J.
1
Expert's answer
2018-03-12T09:50:20-0400
Solution:
2Cl2e= Cl2; n = 2
F – 96500 Q
1 mole of Cl2 (70.9 g) require 2 x 96500 Q of electricity
1.5 x 1012 g require x
X = 2 x 96500 Q x 1.5 x 10^12 g / 70.9 g = 4.08 x 10^15 Q
Per 1 hour: Q = It; I = Q / t = 4.08 x 1015 Q / 3600 s = 1.13 x 10^12 A
The power: P = IU = 1.13 x 1012 A x 4.5 V = 5.1 x 10^12 W = 5.1 x 10^9 kW.

For 1 year process: 365 days = 8760 hours
5.1 x 10^9
kW x 8760 h = 4.47 x 10^16 kW
Answer: 4.47 x 10^16 kW.